pandasurvive:

Two cooks purchased plates and cups at the same store. The first cook bought 4 plates and 5 cups for a total of $25.25. The second cook bought 7 plates and 3 cups for a total of $25.50. What is the price of a plate?

pandasurvive:

I am unsure on how to solve this question.

Hero:

Let p represent the price of one plate and c represent the price of one cup Then the total price of cups and plates purchased by the 1st cook is represented by 4p + 5c = 25.25 And the total price of cups and plates purchased by the 2nd cook is represented by 7p + 3c = 25.50 In other words we have to solve the following system for p: 4p + 5c = 25.25 7p + 3c = 25.50

pandasurvive:

The amount of p and c has to be the same in both equations right?

Hero:

I'm pretty confident you already know the answer to that question.

pandasurvive:

Just making sure.

pandasurvive:

Hero is there a faster way to do it instead of choosing random numbers for p then trying to see if they work?

Vocaloid:

you can use elimination. try multiplying the first equation by 3 and the second equation by 5 and let me know what you get

pandasurvive:

the entire equation? like 4p+5c=25.25 (5)

Vocaloid:

first equation *3 second equation * 5, so: [ 4p+5c=25.25 ] * (3) [ 7p + 3c = 25.50 ] * (5) now try distributing the 3 and 5

pandasurvive:

12p+15c=75.75

pandasurvive:

35p+15c=127.5

Vocaloid:

good, now subtract equation 2 - equation 1 35p - 12p = ? 15c - 15c = ? 127.5 - 75.75 = ?

pandasurvive:

23p=51.75

Vocaloid:

good, so p = ?

pandasurvive:

2.25

pandasurvive:

:o thanks cx

Vocaloid:

good, so plate = $2.25 = your answer